OpenStudy (anonymous):

Find the quotient: 2/3 divided by 4/15 =?

OpenStudy (anonymous):

do you know how to divide fractions ?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

take the fraction that you are dividing by, and " flip " it, then just multiply. Like this... 2/3 divided by 4/15 (flip what you are dividing by, then multiply) 2/3 * 15/4 = 30/12 = 15/6 or 2 1/2 understand ?

OpenStudy (anonymous):

yes. thats what i was messing up on.

OpenStudy (anonymous):

do you understand it now ?

OpenStudy (anonymous):

yes i do.. thank you!

OpenStudy (anonymous):

anytime....thats what I am here for :)